# Interactive Public Transport Infrastructure Analysis through Mobility Profiles: Making the Mobility Transition Transparent

### v1

@@ 6 Conclusion
We have introduced a novel agent-based simulation approach for analyzing and optimizing public transport infrastructure at housing-level granularity. By leveraging housing locations, transport networks, and schedules combined with agent-based simulations, controlled through rapid iterative refinements, our system provides detailed insights into the accessibility and connectivity of public transport networks for members of the public. This approach not only bridges the gap between theoretical planning and practical implementation but also identifies inherent structural issues, thereby highlighting specific areas where interventions are most needed, thus facilitating more informed urban planning decisions. Our findings through a case study in Konstanz, as well as a formative user study, indicate that our approach can effectively identify critical gaps in public transport services and suggest enhancements that are both practical, focusing on economically viable iterative improvements rather than comprehensive overhauls. By enabling detailed and scalable analyses, our system assists policymakers and the interested public, but potentially also city planners, in making data-informed decisions that foster a more equitable and efficient mobility transition. Looking forward, we see substantial opportunities for extending our approach to include (inter-) national data, which could further refine the accuracy and timeliness of transport planning across regional boundaries. By providing both comprehensive insights and support the generation of actionable paths to improvement, our contribution paves the way for more strategic, data-driven urban transport planning and systems, adapting to changing urban landscapes and mobility demands, ultimately contributing to more sustainable cities and improved public infrastructure.

### v2

@@ 6 Conclusion
We have introduced a novel agent-based simulation approach for analyzing and optimizing public transport infrastructure at housing-level granularity. By leveraging housing locations, transport networks, and schedules combined with agent-based simulations, controlled through rapid iterative refinements, our system provides detailed insights into the accessibility and connectivity of public transport networks for members of the public. This approach not only bridges the gap between theoretical planning and practical implementation but also identifies inherent structural issues, thereby highlighting specific areas where interventions are most needed, thus facilitating more informed urban planning decisions. Our findings through a case study and a initial user study, indicate that our approach can effectively identify critical gaps in public transport services and suggest enhancements that are both practical, focusing on economically viable improvements rather than comprehensive overhauls. By enabling detailed and scalable analyses, our system assists policymakers and the interested public, but potentially also city planners, in making data-informed decisions that foster a more equitable and efficient mobility transition. Looking forward, we see substantial opportunities for extending our approach to include (inter-) national data, which could further refine the accuracy and timeliness of transport planning across regional boundaries. By providing both comprehensive insights and support the generation of actionable paths to improvement, our contribution paves the way for more strategic, data-driven urban transport planning and systems, adapting to changing urban landscapes and mobility demands, ultimately contributing to more sustainable cities and improved public infrastructure.
Acknowledgements.
This work has been received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2117 – 422037984.